OpenStudy (anonymous):

Differentiate. y= sec 0 tan 0

OpenStudy (anonymous):

\[y=\sec \theta \tan \theta \]

OpenStudy (anonymous):

use the product rule...

OpenStudy (anonymous):

\[\text{if }y=u \cdot v \text{ then }y\,'=u'\cdot v + u \cdot v\,'\]

OpenStudy (anonymous):

I keep getting the wrong answer

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

\[\sec \theta (\tan ^{2}\theta) + \sec ^{2} (\theta)\]

OpenStudy (anonymous):

\[\left[\sec \left( \theta \right)\tan\left( \theta \right)\right]\tan \left( \theta \right)+\sec \left( \theta \right)\left[ \sec^2\left( \theta \right) \right]=\sec \left( \theta \right)\tan^2\left( \theta \right)+\sec^3\left( \theta \right)\]